02 September 2009 (First night of LBC post-shutdown checkout)

02:17 MST LBCs up and running. Note that during the afternoon HK blue readings had dropped out. So I power cycled HK, but then when I tried to turn on Blue/Red cameras, Blue hinibit counter was too high and LBCB did not initialize. Turned off/on systems and OK now.

02:20 Started series of biases. These look better than last night - banding and noise are not apparent.

02:55 Could not run IDL from lbtmu41,42,43 (IT # 2310). Called Norm who called Stephen Hooper. IDL license server machine is in Tucson not on mountain. Using lbtmu06 to run IDL/dofpia.

03:30 dofpia - converging well on Red - Blue was out of focus and had obstruction. See image at bluepupil.jpg John M. saw that a very thin cardboard or masonite gasket (not the baffle) that is between the cover & the hub was not secured well by the bolts - he secured it.

04:00 Reopening.

04:30 Ran dofpia - red converged quickly, but blue took a few more iterations. Now both are nearly converged and FPIA estimated seeing is about 0.75" (R) and 1.1" (B). Shift "S" out of dofpia because it is trying to meet 400nm limits and not quite there. Want to get on with test of rotator trajectory.

04:36 Had to shut down due to threatening clouds. Had just checked pointing (had to adjust CA by about 50" , however IE was only off by <2") and was editing in coordinates for the TestRotatorTrajectory OB when we shut.

05:30 Reopened. Loaded OB which should have moved to F972 filter, but we had a filter HW failure. Ran dofpia which put in the r-SLOAN filter and this worked.

05:37 Recollimated and running TestRotatorTrajectory0_B31 OB to take 3 dithered 10-sec images at U and F972 --- Red PA = 55 and blue PA = 0. Want to check images to see if trajectory is reasonable.

06:54 The 10-sec exposures showed distorted star images, but hard to tell. Increased exposure time to 50 sec. At this exposure time, we saw jumps and arcs. These images were supposedly guided (but blue was not - did not find guide star - U-BESSEL). Called Norm. I tried unticking "Guiding" in the UI to turn off guiding. This time we see smooth rotational trails. Called Michelle and sent email.

07:00 Repeating experiment - refocussed/collimated and going to repeat OB.

07:20 Repeated OB. Guided show jumps & arcs which are not about center - consistent with what you would expect given the guide star position off the science array, and unguided images show smooth arcs. See, e.g. 070149 (R) and 070149 (B) for guided 50-sec image and 070855 (R) and 070853 (B) for unguided 50-sec image.

07:32 Refocussed/collimated at meridian - RA=23, dec=60deg in preparation for taking 2 superfoc sequences to measure tilt. Note the trailing may affect images, but taking data nonetheless. Seeing is very good - 0.9" B and 0.5" R

07:34 073401... PA=0 and Rotator angle = -180.

07:37 073737... PA=-1.49 and Rot angle = -1.5

LBC> hsel /Repository/lbcb.20090902.073*.fits[1] $I,focusoff,filter
/Repository/lbcb.20090902.073220.fits[1]        -0.080  V-BESSEL
/Repository/lbcb.20090902.073259.fits[1]        -0.040  V-BESSEL
/Repository/lbcb.20090902.073337.fits[1]        0.000   V-BESSEL
/Repository/lbcb.20090902.073416.fits[1]        0.040   V-BESSEL
/Repository/lbcb.20090902.073454.fits[1]        0.080   V-BESSEL
/Repository/lbcb.20090902.073737.fits[1]        -0.080  V-BESSEL
/Repository/lbcb.20090902.073815.fits[1]        -0.040  V-BESSEL
/Repository/lbcb.20090902.073852.fits[1]        0.000   V-BESSEL
/Repository/lbcb.20090902.073930.fits[1]        0.040   V-BESSEL
/Repository/lbcb.20090902.074008.fits[1]        0.080   V-BESSEL
LBC> hsel /Repository/lbcr.20090902.073*.fits[1] $I,focusoff,filter
/Repository/lbcr.20090902.073218.fits[1]        -0.023  r-SLOAN
/Repository/lbcr.20090902.073252.fits[1]        0.017   r-SLOAN
/Repository/lbcr.20090902.073326.fits[1]        0.057   r-SLOAN
/Repository/lbcr.20090902.073401.fits[1]        0.097   r-SLOAN
/Repository/lbcr.20090902.073436.fits[1]        0.137   r-SLOAN
/Repository/lbcr.20090902.073735.fits[1]        -0.023  r-SLOAN
/Repository/lbcr.20090902.073809.fits[1]        0.017   r-SLOAN
/Repository/lbcr.20090902.073843.fits[1]        0.057   r-SLOAN
/Repository/lbcr.20090902.073918.fits[1]        0.097   r-SLOAN
/Repository/lbcr.20090902.073953.fits[1]        0.137   r-SLOAN

08:00 Took an extra/intra focal pair after collimating.

08:08 Recollimated/focussed and taking a Coarse superfoc sequence (step size is 0.08mm instead of 0.04mm) at PA=0, PA=180.

08:23 Recollimated/focussed and taking now a LongCoarse superfoc sequence (step size 0.08mm and 8 vs 5 points) PA=0. Note that images are only 5 sec long, yet LBC log reports that it got a 350 sec tracking trajectory. I thought that for short images no trajectory was called for?
2009/09/02 08:27:18.714218 N B                   got 350 seconds long rotator tracking trajectory (includes telescope preset time)
2009/09/02 08:27:18.714321 N B                   rotator tracking trajectory: "2455076.85229168,0.413747109988377;2455076.85230325,0.413871089878764;2455076.85231483,0.41399506433182;2455076.8523264,0.414119022568801;2455076.85233797,0.414242986131191;2455076.85234955,0.414366937049334;2455076.85236112,0.414490889719806;2455076.85237..."

08:34 Refocus/collim and take another extra/intra focal pair: 083454(B),083452(R) and 083543(B),083541(R) On extra focal Blue pupil images I see bright arc along the rim which is field dependent. On Red I see evidence of residual Z22 - bright annulus in between outer diameter and inner hole. Did not check the intra focal.

08:43 Refocus/collimn and run another typical OB, this time ngc7469guidetest in /home/LBTO/TestOBs - 120 sec images, 5 dithers. Another test of rotator - guiding is on. 085628 (R) shows the arc/jumps that we saw earlier. z-SLOAN guide stars are donuts (on thumbnail). Stopped OB, but this caused LBC to shutdown... I stopped OBs in slew, but does stopping OB during LBC image taking/readout cause a problem? LBC says it is restarting after a crash (09:02). Then an RPC Client/Host communication error for Red camera on startup. Restart... OK now.

09:14 Take a coarse superfoc with B and z. Sequences start with 092627(R) and 092629(B)

09:30 Calling it a night. Powering off LBC - camera, filter, rotator, trackers and HouseKeeping. Shutting down archive machine (lbcarchive2).

problems:
  • rotator trajectory
  • LBC stops ungracefully --- what really happened here? a one-off?
  • z-SLOAN focus offset real or an artifact of poor focus/collimation?
  • need to check collimation table, in particular X, Y limits, but seems decent from object to object. Blue started further out of focus/collimation than Red.But neither required manual input
  • IDL license server machine not on mountain.

OlgaKuhn - 02 Sep 2009
Topic revision: r2 - 02 Sep 2009, OlgaKuhn
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback